OpenStudy (anonymous):

Maya had $27. She spent all the money on buying 3 burgers for $x each and 2 sandwiches for $y each. If Maya had bought 2 burgers and 1 sandwich, she would have been left with $11.

OpenStudy (anonymous):

A student concluded that the price of each burger is $5 and the price of each sandwich is $6. Which statement best justifies whether the student's conclusion is correct or incorrect? The student's conclusion is correct because the solution to the system of equations 3x + 2y = 11 and 2x + y = 16 is (5, 6). The student's conclusion is incorrect because the solution to the system of equations 3x – 2y = 11 and 2x – y = 16 is (5, 6). The student's conclusion is correct because the solution to the system of equations 3x + 2y = 27 and 2x + y = 16 is (5, 6). The student's conclusion is incorrect because the solution to the system of equations 2x + 3y = 27 and x + 2y = 16 is (5, 6).

OpenStudy (anonymous):

burgers sound so good right now omg im so sorry

OpenStudy (anonymous):

... lol

OpenStudy (anonymous):

do you think you can help me?

OpenStudy (anonymous):

yea I can try

OpenStudy (anonymous):

okay thx

OpenStudy (anonymous):

its definitely not B because in the problem it says that the student concluded the numbers 5 and 6, so it couldn't be incorrect

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

im going to go with C.........

OpenStudy (anonymous):

okay can you tell me why?

OpenStudy (anonymous):

I cant D: im sorry!

OpenStudy (anonymous):

okay